# Seungwon Baek b,c    P. Ko d    Yuji Omura e    and Chaehyun Yu

### v1

@@ Abstract
In this paper, we study two Higgs doublet models with gauged $U(1)_{H}$ symmetry, motivated by the excesses around 96 GeV reported by the CMS collaboration in the searches for light resonances decaying to two photons and two $\tau$’s. In this model, one Higgs doublet field is charged under the $U(1)_{H}$ symmetry to avoid tree-level flavor changing neutral currents. The extra gauge symmetry requires extra chiral fermions, to satisfy the anomaly-free conditions. We analyze the signals of the light resonances, taking into account the contribution of the extra fermions, and discuss the consistency with the experimental results in this model.
@@ 1 Introduction
The Standard Model (SM) has been established as a theory describing particle physics. Almost all predictions of the SM are consistent with experimental results and the Higgs particle was finally discovered at the LHC ATLAS:2012yve ; CMS:2012qbp . It is certain that there are still large uncertainties in some observables, so that new physics may exist in the energy region that can be explored at the LHC. The new model, however, should not drastically modify the SM predictions. A lot of candidates for new physics have been proposed, motivated by problems in the SM. The mysteries, especially, concerned with the origin of the electroweak (EW) scale and the vacuum structure of our universe seems to suggest new particles around the EW scale. Thus, some extensions of the SM model that reveals new aspects of our universe may be confirmed near future.
@@
If those deviations are originated from new resonances that reside around 96 GeV, one of the good candidates is a neutral scalar originated from scalar fields that contribute to the EW symmetry breaking Biekotter:2023oen ; Biekotter:2019kde ; Biekotter:2019mib ; Biekotter:2020cjs ; Biekotter:2021qbc ; Heinemeyer:2021msz ; Biekotter:2022jyr ; Biekotter:2023jld ; Bhattacharya:2023lmu ; Banik:2023ntf . The possibility that extra scalar fields may exist has been discussed in many works. Adding extra scalar fields is surely one simple way to extend the SM without disturbing the anomaly-free conditions. Scalar fields may play a role in breaking extra gauge symmetries. In fact, extra neutral scalars that are related to gauge symmetry breaking and reside around 96 GeV have been proposed motivated by the excesses, and the signals have been studied in many works Kundu:2019nqo ; Cao:2019ofo ; Abdelalim:2020xfk ; Belyaev:2023xnv ; Aguilar-Saavedra:2023tql ; Dutta:2023cig ; Ellwanger:2023zjc ; Cao:2023gkc ; Arcadi:2023smv ; Mulaudzi:2023khg ; Liu:2024cbr ; Cao:2024axg ; Kalinowski:2024uxe ; Ellwanger:2024txc ; Ellwanger:2024vvs ; Arhrib:2024wjj ; Benbrik:2024ptw ; Lian:2024smg ; Khanna:2024bah ; BrahimAit-Ouazghour:2024img ; Ashanujjaman:2023etj ; Banik:2023vxa ; Coloretti:2023yyq ; Dong:2024ipo ; Borah:2023hqw ; Ahriche:2023hho ; Ahriche:2023wkj ; Chen:2023bqr ; Dev:2023kzu ; Wang:2024bkg ; YaserAyazi:2024hpj .
@@
The Higgs-mediated FCNC problem in generic 2HDMs can be cured a la by Natural Flavor Conservation Criterion by Glashow and Weinberg Glashow:1976nt . A simple realization of their criterion is to introduce $Z_{2}$ symmetry under which two Higgs doublets are different $Z_{2}$ parity. Then the SM chiral fermions will also carry different $Z_{2}$ charges. There are four different charge assignments for which Yukawa couplings of the SM fermions are allowed in such a way that the fermions of the same electric charge gets their masses from one and the same Higgs doublet, either $H_{1}$ or $H_{2}$, but not from both. Namely 2HDMs that do not predict tree level FCNCs are classified into four types. In the Type-I 2HDM, only one Higgs field is coupled to the SM fermions. In the Type-II 2HDM, one Higgs field ($H_{1}$) couples with right-handed down-type quarks and charged leptons, and the other Higgs field ($H_{2}$) couples with right-handed up-type quarks and neutrinos. It is possible to consider a model where the couplings are flipped: $H_{1}$ couples with right-handed down-type quarks and neutrinos, and the $H_{2}$ couples with right-handed up-type quarks and charged leptons. We call this model the Type-Y 2HDM. In the fourth 2HDM, $H_{1}$ couples with quarks and $H_{2}$ couples with leptons. Then this $Z_{2}$ symmetry is assumed to be softly broken by dim-2 operator, $m_{12}^{2}H_{1}^{\dagger}H_{2}+H.c.$, in order to increase the masses of the scalars originated from $H_{1}$ and $H_{2}$.
In this paper, we consider gauged $U(1)_{H}$ symmetry instead of softly broken $Z_{2}$ symmetry, by assigning a non-vanishing $U(1)_{H}$ charge to $H_{2}$ Ko:2012hd ; Ko:2013zsa ; Ko:2014uka ; Ko:2015fxa . In addition, extra complex singlet scalar field, $\Phi$, is also introduced in order to break $U(1)_{H}$ spontaneously. As discussed in Ref. Ko:2012hd , there are several possible setups in 2HDMs with gauged $U(1)_{H}$. In this work, we concentrate on the Type-II 2HDM and the Type-Y 2HDM motivated by the excesses around 96 GeV. In Sec. 2, we discuss the setups of our 2HDMs. In Sec. 3, we summarize experimental constraints relevant to our models. In Sec. 4, the analyses concerned with the excesses are shown. Sec. 5 is devoted to summary. In Appendix A, the constraints from the condition for the scalar potential to be bounded from below is introduced. In Appendix B, the relevant effective couplings are summarized.
@@ 2 The 2HDM with gauged $U(1)_{H}$ symmetry
We consider 2HDMs with extra $U(1)_{H}$ gauge symmetry, where one of the Higgs field $H_{2}$ is charged under the $U(1)_{H}$ gauge symmetry, while the other $H_{1}$ is not charged. Some SM fermions may be also charged under $U(1)_{H}$ so that phenomenologically viable Yukawa couplings for the SM fermions are allowed by the presumed gauge symmetry. Below, we discuss the detail based on Refs. Ko:2012hd ; Ko:2013zsa ; Ko:2014uka ; Ko:2015fxa .
@@
$U(1)_{Y}$ $1/2$ $1/2$ 00
$U(1)_{H}$ 00 $1$ $-1$
Table 1: The charge assignment to the scalar fields.
@@
where $k=1,2$. We emphasize that, compared to the usual 2HDMs, two operators are missing in our model due to the $U(1)_{H}$ gauge symmetry: the soft $Z_{2}$ breaking dim-2 operator, $m_{12}^{2}H_{1}^{\dagger}H_{2}$, and and the $\lambda_{5}(H_{1}^{\dagger}H_{2})^{2}+h.c.$ term Still the $H_{1}^{\dagger}H_{2}+h.c.$ term can be realized from the $H_{1}^{\dagger}H_{2}\Phi+h.c$ terms after the singlet field $\Phi$ develops a non-vanishing VEV.
@@
After the electroweak (EW) and $U(1)_{H}$ gauge symmetry breaking, three CP-even neutral scalar fields, $h_{i}\equiv(h_{1},h_{2},h_{\Phi})$, mix among themselves, and the $3\times 3$ mass matrix is diagonalized by three physical states, $S_{i}\equiv(\tilde{h},h,H)$. $h$ is identified as the Higgs boson with the mass of 125 GeV while $\tilde{h}$ is as the $96$ GeV scalar boson. The remaining $H$ is an additional scalar boson. The mixing between three scalar bosons are defined by
@@
In this paper, we do not pay attention to the extra fermions too much because they strongly depends on the model construction which is not unique. Our motivation is to study the effects of the extra fermions to the scalar boson decays such as $\tilde{h}\rightarrow\gamma\gamma$ through the loops. Thus the extra fermion contributions to the loop diagrams must be considered collectively, that is the sum of all the fermions in the model, which is quite cumbersome. Instead, we assume a vectorlike charged lepton and quark as a representative of collection of the extra fermions in the loop, denoting the Yukawa couplings of extra quarks and leptons as $y_{D}$ and $y_{L}$, respectively. In our phenomenological analysis, those parameters are fixed at $y_{D}=y_{L}=1$.
@@ 3.1 Theoretical constraints
First, we consider the constraints from the vacuum stability conditions for nonzero VEVs. The requirement for the scalar potential to be stable and bounded from below constrain the dimensionless couplings in the scalar potential. Following the approach in Ref. Arhrib:2011uy , we find the constraints on the dimensionless couplings, which are presented in appendix A.
@@
00 0 $=$ $x^{3}-4(3\lambda_{1}+3\lambda_{2}+2\lambda_{\Phi})x^{2}$
@@ 3.2 Electroweak precision observables
In order to study the physical effects on the electroweak precision observables (EWPOs) in this model, full calculations of the relevant amplitudes at the one-loop level is required, which is quite involved. Instead we calculate the $\Delta T$ parameter approximately. Due to the $Z$-$Z^{\prime}$ mixing, the mass of the $Z$ boson is shifted
@@
for $x\neq y$ while $F(x,y)=0$ for $x=y$. The effective couplings, $g_{W^{\pm}H^{\mp}A}$ and so on, are defined in Appendix B. This loop contribution to $T$ in the case of 2HDMs with $Z_{2}$ symmetry vanishes. However, in our model, the EW symmetry breaking makes the $Z^{\prime}$ boson massive by absorbing one of degrees of freedom in the scalar fields. Therefore, the presence of the $U(1)_{H}$ gauge symmetry implies that the contribution from the $Z$-$Z^{\prime}$ mixing must be taken into consideration together with the scalar loop corrections. Finally, the sum of two contributions leads to
@@
where the measured value is ParticleDataGroup:2022pth
$T^{\rm exp}=0.09\pm 0.07.$ (33)
@@ 3.3 Flavor physics
The charged Higgs boson mass $m_{H^{\pm}}$ and $\tan\beta$ are strongly constrained by flavor physics, in particular, the $b\to s\gamma$ decay. The constraints on the 2HDMs from flavor physics including the experimental results at the LHC were analyzed in Ref. Arbey:2017gmh . The constrains on the charged Higgs boson mass and $\tan\beta$ are the same as those in the 2HDMs. Thus, we take the lower bound of the charged Higgs boson mass to be 600 GeV and set the bound on $\tan\beta$ from 1 to 3 depending on the charged Higgs boson mass.
@@ 3.4 Experiments for scalar bosons
There have been a lot of searches for new scalar bosons at the LEP, Tevatron, and LHC, and lots of data have been accumulated. However, there is no definite signal for a new scalar boson so far. These search data except for the resonance data around $96$ GeV will strongly constrain our model. We apply these experimental constraints for the additional scalar bosons using the public code HiggsBounds Bechtle:2020pkv where the exclusion limit is 95% C.L..
In our model, the 125 GeV scalar boson will mix with other neutral scalar bosons. Therefore, this mixings change the predictions for the properties of the observed Higgs boson. Then the precision observables for productions and decays of the 125 GeV Higgs boson also constrain our model strongly. We use the public code HiggsSignals Bechtle:2020uwn to test the 125 GeV Higgs boson in our model. The code performs a $\chi^{2}$ test for signal and decay rates of the 125 GeV Higgs boson against the prediction in the SM. Then we put the condition that the p-value evaluated from HiggsSignals is less than $0.05$.

### v2

@@ Abstract
In this paper, we study two Higgs doublet models with gauged $U(1)_{H}$ symmetry instead of the usual softly broken $Z_{2}$ symmetry, motivated by the excesses around 96 GeV reported by the CMS collaboration in the searches for light resonances decaying to two photons and two $\tau$’s. In this model, one Higgs doublet field is charged under the $U(1)_{H}$ symmetry, thereby one can avoid tree-level flavor changing neutral currents. The extra $U(1)_{H}$ gauge symmetry requires extra chiral fermions, in order to satisfy the anomaly-free conditions. We analyze the signals of the light resonances, taking into account the contribution of the extra fermions, and discuss the consistency with the experimental results in this model.
@@ 1 Introduction
The Standard Model (SM) has been established as a theory describing particle physics. Almost all predictions of the SM are consistent with experimental results and the Higgs particle was finally discovered at the LHC ATLAS:2012yve ; CMS:2012qbp . It is certain that there are still large uncertainties in some observables, so that new physics may exist in the energy region that can be explored at the LHC. The new model, however, should not drastically modify the SM predictions. A lot of candidates for new physics have been proposed, motivated by problems in the SM. The mysteries, especially, concerned with the origin of the electroweak (EW) scale and the vacuum structure of our universe seem to suggest new particles around the EW scale. Thus, some extensions of the SM model that reveal new aspects of our universe may be confirmed near future.
@@
If those deviations are originated from new resonances that reside around 96 GeV, one of the good candidates is a neutral scalar originated from scalar fields that contribute to the EW symmetry breaking Biekotter:2023oen ; Biekotter:2019kde ; Biekotter:2019mib ; Biekotter:2020cjs ; Biekotter:2021qbc ; Heinemeyer:2021msz ; Biekotter:2022jyr ; Biekotter:2023jld ; Bhattacharya:2023lmu ; Banik:2023ntf . The possibility that extra scalar fields may exist has been discussed in many works. Adding extra scalar fields is surely one simple way to extend the SM without disturbing the anomaly-free conditions. Scalar fields may play a role in breaking extra gauge symmetries. In fact, extra neutral scalars that are related to gauge symmetry breaking and reside around 96 GeV have been proposed motivated by the excesses, and the signals have been studied in many works Kundu:2019nqo ; Cao:2019ofo ; Abdelalim:2020xfk ; Belyaev:2023xnv ; Aguilar-Saavedra:2023tql ; Dutta:2023cig ; Ellwanger:2023zjc ; Cao:2023gkc ; Arcadi:2023smv ; Mulaudzi:2023khg ; Liu:2024cbr ; Cao:2024axg ; Kalinowski:2024uxe ; Ellwanger:2024txc ; Ellwanger:2024vvs ; Arhrib:2024wjj ; Benbrik:2024ptw ; Lian:2024smg ; Khanna:2024bah ; BrahimAit-Ouazghour:2024img ; Ashanujjaman:2023etj ; Banik:2023vxa ; Coloretti:2023yyq ; Dong:2024ipo ; Borah:2023hqw ; Ahriche:2023hho ; Ahriche:2023wkj ; Chen:2023bqr ; Dev:2023kzu ; Wang:2024bkg ; YaserAyazi:2024hpj ; Gao:2024qag ; Haisch:2017gql ; Fox:2017uwr ; Biekotter:2021ovi ; Azevedo:2023zkg .
@@
The Higgs-mediated FCNC problem in generic 2HDMs can be cured a la by Natural Flavor Conservation Criterion by Glashow and Weinberg Glashow:1976nt . A simple realization of their criterion is to introduce $Z_{2}$ symmetry under which two Higgs doublets has different $Z_{2}$ parity. Then the SM chiral fermions will also carry different $Z_{2}$ charges. There are four different charge assignments for which Yukawa couplings of the SM fermions are allowed in such a way that the fermions of the same electric charge get their masses from one and the same Higgs doublet, either $H_{1}$ or $H_{2}$, but not from both. Namely 2HDMs that do not predict tree level FCNCs are classified into four types. In the Type-I 2HDM, only one Higgs field is coupled to the SM fermions. In the Type-II 2HDM, one Higgs field ($H_{1}$) couples with right-handed down-type quarks and charged leptons, and the other Higgs field ($H_{2}$) couples with right-handed up-type quarks and neutrinos. It is possible to consider a model where the couplings are flipped: $H_{1}$ couples with right-handed down-type quarks and neutrinos, and the $H_{2}$ couples with right-handed up-type quarks and charged leptons. We call this model the Type-Y 2HDM. In the fourth 2HDM, $H_{1}$ couples with quarks and $H_{2}$ couples with leptons. Then this $Z_{2}$ symmetry is assumed to be softly broken by dim-2 operator, $m_{12}^{2}H_{1}^{\dagger}H_{2}+H.c.$, in order to increase the masses of the scalars originated from $H_{1}$ and $H_{2}$.
In this paper, we consider gauged $U(1)_{H}$ symmetry instead of softly broken $Z_{2}$ symmetry, by assigning a non-vanishing $U(1)_{H}$ charge to $H_{2}$ Ko:2012hd ; Ko:2013zsa ; Ko:2014uka ; Ko:2015fxa . In addition, extra complex singlet scalar field, $\Phi$, is also introduced in order to break $U(1)_{H}$ spontaneously. As discussed in Ref. Ko:2012hd , there are several possible setups in 2HDMs with gauged $U(1)_{H}$. In this work, we concentrate on the Type-II 2HDM and the Type-Y 2HDM motivated by the excesses around 96 GeV. In Sect. 2, we discuss the setups of our 2HDMs. In Sect. 3, we summarize experimental constraints relevant to our models. In Sect. 4, the analyses concerned with the excesses are shown. Section 5 is devoted to summary. In Appendix A, the constraints from the condition for the scalar potential to be bounded from below is introduced. In Appendix B, the relevant effective couplings are summarized.
@@ 2 The 2HDM with gauged $U(1)_{H}$ symmetry
We consider 2HDMs with extra $U(1)_{H}$ gauge symmetry, where one of the Higgs fields, $H_{2}$, is charged under the $U(1)_{H}$ gauge symmetry, while the other $H_{1}$ is not charged. Some SM fermions may be also charged under $U(1)_{H}$ so that phenomenologically viable Yukawa couplings for the SM fermions are allowed by the presumed gauge symmetry. Below, we discuss the detail based on Refs. Ko:2012hd ; Ko:2013zsa ; Ko:2014uka ; Ko:2015fxa .
@@
$U(1)_{Y}$ $1/2$ $1/2$ 0
$U(1)_{H}$ 0 11 1-1
Table 1: The charge assignment to the scalar fields. Irreducible representations of $SU(3)_{c}$ and $SU(2)$ are denoted by their dimensions in boldface.
@@
where $k=1,2$. We emphasize that, compared to the usual 2HDMs, two operators are missing in our model due to the $U(1)_{H}$ gauge symmetry: the soft $Z_{2}$ breaking dim-2 operator, $m_{12}^{2}H_{1}^{\dagger}H_{2}$, and the $\lambda_{5}(H_{1}^{\dagger}H_{2})^{2}+h.c.$ term. Still the $H_{1}^{\dagger}H_{2}+h.c.$ term can be realized from the $H_{1}^{\dagger}H_{2}\Phi+h.c$ terms after the singlet field $\Phi$ develops a non-vanishing VEV.
@@
After the electroweak (EW) and $U(1)_{H}$ gauge symmetry breaking, three CP-even neutral scalar fields, $h_{i}\equiv(h_{1},h_{2},h_{\Phi})$, mix among themselves, and the $3\times 3$ mass matrix is diagonalized by three physical states, $S_{i}\equiv(\tilde{h},h,H)$. $h$ is identified as the Higgs boson with the mass of 125 GeV while $\tilde{h}$ is as the $96$ GeV scalar boson. The remaining $H$ is an additional scalar boson. The mixing between three scalar bosons is defined by
@@
In this paper, we do not pay attention to the extra fermions too much because they strongly depend on the model construction which is not unique. Our motivation is to study the effects of the extra fermions to the scalar boson decays such as $\tilde{h}\rightarrow\gamma\gamma$ through the loops. Thus, the extra fermion contributions to the loop diagrams must be considered collectively, that is the sum of all the fermions in the model, which is quite cumbersome. Instead, we assume a vectorlike charged lepton and quark as a representative of collection of the extra fermions in the loop, denoting the Yukawa couplings of extra quarks and leptons as $y_{D}$ and $y_{L}$, respectively. In our phenomenological analysis, those parameters are fixed at $y_{D}=y_{L}=1$.
@@ 3.1 Theoretical constraints
First, we consider the constraints from the vacuum stability conditions for nonzero VEVs. The requirement for the scalar potential to be stable and bounded from below constrains the dimensionless couplings in the scalar potential. Following the approach in Ref. Arhrib:2011uy , we find the constraints on the dimensionless couplings, which are presented in appendix A.
@@
$0$ $=$ $x^{3}-4(3\lambda_{1}+3\lambda_{2}+2\lambda_{\Phi})x^{2}$
@@ 3.2 Electroweak precision observables
In order to study the physical effects on the electroweak precision observables (EWPOs) in this model, full calculations of the relevant amplitudes at the one-loop level are required, which are quite involved. Instead, we calculate the $\Delta T$ parameter approximately. Due to the $ZZ^{\prime}$ mixing, the mass of the $Z$ boson is shifted
@@
for $x\neq y$ while $F(x,y)=0$ for $x=y$. The effective couplings, $g_{W^{\pm}H^{\mp}A}$ and so on, are defined in Appendix B. This loop contribution to $T$ in the case of 2HDMs with $Z_{2}$ symmetry vanishes. However, in our model, the EW symmetry breaking makes the $Z^{\prime}$ boson massive by absorbing one of degrees of freedom in the scalar fields. Therefore, the presence of the $U(1)_{H}$ gauge symmetry implies that the contribution from the $ZZ^{\prime}$ mixing must be taken into consideration together with the scalar loop corrections. Finally, the sum of two contributions leads to
@@
where the measured value is ParticleDataGroup:2024cfk
$T^{\rm exp}=0.04\pm 0.03.$ (33)
@@ 3.3 Flavor physics
The charged Higgs boson mass $m_{H^{\pm}}$ and $\tan\beta$ are strongly constrained by flavor physics, in particular, by the $b\to s\gamma$ decay. The constraints on the 2HDMs from flavor physics including the experimental results at the LHC were analyzed in Ref. Arbey:2017gmh ; Haller:2018nnx . In the Type-II 2HDM, the lower bound of the charged Higgs mass is about 600 GeV, and this is almost independent of $\tan\beta$. The study of $B\to X_{s}\gamma$ in Ref. Misiak:2017bgg also has led the same lower bound. In Ref. Misiak:2020vlo , the authors improve the SM prediction of $B\to X_{s}\gamma$: $Br(B\to X_{s}\gamma)=(3.40\pm 0.17)\times 10^{-4}$ Misiak:2020vlo . They also comment that the bound of $m_{H^{\pm}}$ becomes very severe: $m_{H^{\pm}}\gtrsim 800$ GeV Misiak:2020vlo . [^1] See also Ref. Li:2024kpd . One of the reasons why the bound becomes so strong is that the improved SM prediction is much higher than the experimental result, $Br(B\to X_{s}\gamma)=(3.32\pm 0.15)\times 10^{-4}$ HFLAV:2016hnz . In the Type-II 2HDM, the charged Higgs contribution to $B\to X_{s}\gamma$ interferes with the SM contribution in a constructive manner, so that such new physics contribution is strongly limited.
Recently, the Heavy Flavor Averaging Group (HFLAV) announced the new result, $Br(B\to X_{s}\gamma)=(3.49\pm 0.19)\times 10^{-4}$ HeavyFlavorAveragingGroupHFLAV:2024ctg , that is a little larger than the SM prediction. If the latest HFLAV result is used, the $m_{H^{\pm}}$ bound is expected to be relaxed. We calculate $Br(B\to X_{s}\gamma)$ using SuperIso v4.1 Mahmoudi:2008tp and compare the new HFLAV result. We derive the lower bound of the charged Higgs mass as $m_{H^{\pm}}\gtrsim 500$ GeV, requiring that our prediction is within $2\sigma$ of the new HFLAV result. The more detailed study is beyond our scope, but more accurate analysis of $B\to X_{s}\gamma$ may improve the bound.
Since both bounds are meaningful, we present plots for two cases: $m_{H^{\pm}}>500$ GeV and $m_{H^{\pm}}>800$ GeV. We refer to the former bound as the HFLAV cut and the latter as the tight cut. We note that the bound from $R_{\gamma}=Br_{(s+d)\gamma}/Br_{cl\bar{\nu}}$ is stronger than that from $Br(B\to X_{s}\gamma)$ Li:2024kpd . Therefore, for the tight cut, we apply the constraints on $m_{H^{\pm}}$ and $\tan\beta$ as provided in Ref. Li:2024kpd , where both bounds are simultaneously considered. However, the HFLAV group has not yet reported a value for $R_{\gamma}$. In Ref. Li:2024kpd , the bounds from $Br(B\to X_{s}\gamma)$ and $R_{\gamma}$ differ by about 50 GeV in the lower limit on $m_{H^{\pm}}$. Taking this difference into account, we present plots using the constraint $m_{H^{\pm}}>500$ GeV for the HFLAV cut. In addition, we have tested our models with a more conservative constraint, $m_{H^{\pm}}>550$ GeV, motivated by the bound from $R_{\gamma}$. However, we find that the resulting behavior is not significantly affected by the 50 GeV difference in the $m_{H^{\pm}}$ bound.
Another stringent constraint from flavor physics arises from the $B_{s}\to\mu^{+}\mu^{-}$ decay. In the Type-II 2HDM, this constraint is significant in both the low and high $\tan\beta$ regimes, i.e., $\tan\beta\sim 1$ and $\tan\beta\sim 20$. In contrast, in the Type-Y 2HDM, only the low $\tan\beta$ regime is affected. We apply these constraints to both the HFLAV and tight cuts as provided in Ref. Li:2024kpd . In our model, there are extra quarks that may lead destructive interference with the SM one. The study of flavor physics concerned with the extra quarks would not be simple because the other flavor observables as well as the collider constraints come into play. The detailed study of flavor physics is our future work.
@@ 3.4 Experiments for scalar bosons
There have been a lot of searches for new scalar bosons at the LEP, Tevatron, and LHC, and lots of data have been accumulated. However, there is no definite signal for a new scalar boson so far. These search data except for the resonance data around $96$ GeV will strongly constrain our model. We apply these experimental constraints for the additional scalar bosons using the public code HiggsBounds Bechtle:2020pkv which is encoded in HiggsTools Bahl:2022igd .
In our model, the 125 GeV scalar boson will mix with other neutral scalar bosons. Therefore, these mixings change the predictions for the properties of the observed Higgs boson. Then the precision observables for productions and decays of the 125 GeV Higgs boson also constrain our model strongly. We use the public code HiggsSignals Bechtle:2020uwn to test the 125 GeV Higgs boson in our model. In analysis, we utilize HiggsTools Bahl:2022igd which contains the most recent version of HiggsSignals. In order to ensure that the 125 GeV Higgs boson in our model mimics the behavior of the SM Higgs boson, we impose the conditions $\Delta\chi_{125}^{2}<6.18$ and $\Delta\chi^{2}<0$ Chen:2023bqr , where $\Delta\chi_{125}^{2}$ is the difference of $\chi^{2}$ for the measurement of the 125 GeV Higgs boson between our model and the SM, while $\Delta\chi^{2}$ is that of the combined $\chi^{2}$ of the measurements at the 96 GeV excess as well as those of the 125 GeV Higgs Boson.
3.5 Constraints from $t\bar{t}\tau^{+}\tau^{-}$ production at the LHC
The 96 GeV scalar boson couples to both the top quark and tau lepton. Thus, the production of $t\bar{t}\tau^{+}\tau^{-}$ at the LHC can constrain the scalar boson Iguro:2022dok  [^2] See also Ref. Iguro:2022fel .. Based on the ATLAS Run 2 full data ATLAS:2022yrq , the authors in Ref. Iguro:2022dok have proposed the upper bounds of the Yukawa couplings, $\rho^{tt,\tau\tau}_{\tilde{h}}$, normalized by $\sqrt{2}$. Recently, the CMS collaboration also announced the result of the new physics search in the $t\bar{t}\tau^{+}\tau^{-}$ channel CMS:2024ulc , and the upper bound of the cross section is about $0.6$ times lower than those in Ref.Iguro:2022dok . Using the CMS results, we require that the Yukawa couplings should satisfy
$1.07\times\left|\rho_{\tilde{h}}^{tt}\right|^{2}{\rm Br}\left(\widetilde{h}\to\tau\overline{\tau}\right)$ $<0.03$ , (34)
where $\rho_{\tilde{h}}^{tt}=\frac{g_{\tilde{h}}^{t,\textrm{Y}}m_{t}}{v}$ is defined. It is worthwhile to mention that these constraints are based on the analysis of the tree-level calculation for the $t\bar{t}\tau^{+}\tau^{-}$ production Iguro:2022dok . With QCD corrections to the process, the bounds may be slightly modified.